# Diversity Methods for Improving Convergence and Accuracy of Quantum Error Correction Decoders Through Hardware Emulation

### v1

@@ Abstract
Understanding the impact of accuracy and speed when quantum error correction (QEC) decoders transition from floating-point software implementations to finite-precision hardware architectures is crucial for resource estimation on both classical and quantum sides. The final performance of the hardware implementation influences the code distance, affecting the number of physical qubits needed, and defines connectivity between quantum and classical control units, among other factors like refrigeration systems.
This paper introduces a hardware emulator to evaluate QEC decoders using real hardware instead of software models. The emulator can explore $10^{13}$ different error patterns in 20 days with a single FPGA device running at 150 MHz, guaranteeing the decoder’s performance at logical rates of $10^{-12}$, the requirement for most quantum algorithms. In contrast, an optimized C++ software on an Intel Core i9 with 128 GB RAM would take over a year to achieve similar results. The emulator also enables storing patterns that generate logical errors for offline analysis and to design new decoders.
Using results from the emulator, we propose a diversity-based method combining several belief propagation (BP) decoders with different quantization levels. Individually, these decoders may show subpar error correction, but together they outperform the floating-point version of BP for quantum low-density parity-check (QLDPC) codes like hypergraph or lifted product. Preliminary results with circuit-level noise and bivariate bicycle codes suggest hardware insights can also improve software. Our diversity-based proposal achieves a similar logical error rate as BP with ordered statistics decoding, with average speed improvements ranging from 30% to 80%, and 10% to 120% in worst-case scenarios, while reducing post-processing algorithm activation by 47% to 96.93%, maintaining the same accuracy.
@@ 1 Introduction
The design of decoding algorithms that obtain high accuracy and low latency for quantum low-density parity-check (QLDPC) codes [1] has been a very active area during the last decade, especially over the previous six years. Researchers have been working in low-complexity decoding algorithms to achieve the performance of belief-propagation (BP) combined with ordered-statistics decoding (OSD) [2], [3], which mitigates the effect of degeneracy [4], but minimizing the operational demands of OSD, which introduces scalability challenges, especially when the detector error model is considered [5], [6].
@@
Some of these proposals focus on avoiding OSD-like solutions that require solving systems of linear equations, which implies the computation of matrix inversion through Gaussian elimination. These alternatives seek to enhance efficiency by combining different BP decoders or modifying internal rules for the update of the different nodes involved in the decoding process. Examples of these include Belief Propagation Guided Decimation (BPGD) [7], [8], Stabilizer Inactivation (SI) [9], and Check-Agnosia (CA) [10]. On the other side, some decoding algorithms aim to simplify, rather than eliminate, OSD, such as the Localized Statistics Decoder (BP+LSD) [11] and the Ordered Tanner Forest (BP+OTF) [12]. Regardless of the approach taken, all solutions begin with some variant of BP, leading to further exploration of BP enhancements to improve error correction performance. This includes modifications to BP’s scheduling, such as the random reordering of node updates [13], or introducing additional noise or perturbations to address the error floor problem [14].
@@
However, several important questions remain unanswered regarding these algorithms. First, most quantum algorithms require a logical error rate between $10^{-10}$ and $10^{-13}$ [15]. To ensure that, it is necessary to know the impact of finite-precision architectures on the final error correction performance of the decoders when implemented in hardware, especially in the low logical error rate regions. To obtain statistically significant results, at least $10^{12}$ or $10^{15}$ experiments need to be run, which would be a bottleneck in software. On the other hand, these quantum algorithms require larger-scale quantum devices. Given the scaling challenges, it is reasonable to believe that floating-point operations may not be power or time-efficient, and accuracy should be evaluated with this limitation in mind [16]. Also, studying the potential advantage in the characteristics of the hardware architectures remains uncertain.
@@
Answering these questions requires the implementation of an emulator. Although designing and rigorously verifying such an emulator is time-consuming, it is essential to address with warranties some of the previous problems. The accuracy of the final device will ultimately depend on the hardware architecture rather than the software implementation. Verification in the region of lower logical error rates cannot be realistically achieved through alternative methods, given limited resources. Additionally, the unique aspects of finite precision implementations may lead to the development of new decoders that take advantage of the deviations introduced by these architectures.
@@
Finally, it makes sense to start testing more realistic models of BP, as it underpins all the previously mentioned state-of-the-art decoders. Even if other approaches, such as those based on neural networks and BP, are explored [17], utilizing a hardware emulator will enable more accurate training for achieving logical error rates below $10^{-12}$ (as the samples will be finite precision too), which is currently not feasible with software-based methods.
@@
  • The architecture of a hardware emulator to benchmark QEC real-time decoders and to facilitate offline analysis of the derived information. This emulator enables a significant reduction in the time required to explore low error rates down to $10^{-12}$, allowing this process to be completed in days instead of the years it would take using a software model. Additionally, the analysis of finite-precision BP decoders provides some valuable insights.
  • The design of a BP-based decoder with the knowledge obtained after studying, through the emulator, the diverse levels of noise generated by the quantization schemes. This proposal improves the logical error rate without post-processing. Similarly to the one in [18], in which the authors combined several noisy decoders to generate highly accurate decoding predictions. However, in our proposal, the noisy decoder is BP instead of minimum-weight perfect matching (MWPM), and the noise is not added artificially but is part of the nature of the hardware architectures, which are carefully selected. Finally, due to the nature of BP, it is not necessary to have a degree of consensus among the decoders to increase confidence; we only wait for the one that converges first, following some priorities in between the different quantization schemes. Our solution is based on just four decoders that can share the hardware in groups of two without impact in latency and a minimum memory overhead. The codes under test are hypergraph product and lifted product QLDPC codes [3], [19].
  • The definition of a decoder for circuit level noise that reduces the number of iterations required by BP while at the same time reduces the number of calls for the post-processor obtaining at least the same logical error correction as BP+OSD for bivariate bicycle codes of different lengths and distances [20].
The structure of the paper is detailed as follows: Section II describes the proposed emulator, detailing the components of the architecture and the different configuration parameters. Section III focuses on the verification of the platform by testing the performance of various QLDPC codes with check-node degrees of six and eight, from which some interesting conclusions are drawn. Section IV presents the first diversity approach aimed at improving the accuracy of BP decoders with a moderate overhead in hardware resources. This approach leverages the quantization noise introduced by finite precision architectures to achieve low-latency solutions. Section V introduces a second diversity proposal based on different BP implementations, which can lead to reduced hardware requirements and shorter execution times by minimizing the number of post-processing executions. Finally, Section VI concludes the paper and suggests future research lines.
@@ 2 Proposed Emulator
The architecture depicted in Fig. 1 consists of three main layers: noise and input stimulus generation, an input/output parameter interface, and a communication interface. Additionally, there is a transversal control layer that orchestrates the entire emulator. This structure is adaptable for any type of QEC decoder that utilizes syndromes or detectors as inputs and produces estimated vectors or observables as outputs. Specifically, this work analyzes different versions of BP, as improvements in BP can significantly enhance the performance of the most accurate decoders currently available in the field. It is important to note that the implementation is vendor-agnostic, meaning it can be executed on any FPGA device, and it does not rely on proprietary IP cores.
@@
Next, we detail each of the modules.
2.0.1 Noise and input stimulus generation
@@
At the output of each of these generators, a comparator is implemented. This comparator has one of the inputs wired to a programmable threshold, whose value depends on the error model and the physical error rate. The output of the random generator will indicate if a certain qubit will suffer an X, Y, or Z error.
@@
Depending on the FPGA device and the length of the code under test, $n$, having $NG=n$ random noise generators may consume too many hardware resources, leaving insufficient space for the QEC decoder. The parameter $NG$ is generic and configurable in synthesis time. After the comparators, there is one register that receives $NG$ bits in parallel and outputs $n$. So, when $NG<n$, the number of clock cycles to generate an entire error pattern is $\left\lceil{NG/n}\right\rceil$. When this error sequence is generated, it is stored in a parallel register to start producing the next noise sequence while the decoder is working. This minimizes idle periods.
@@
After the register is updated with a new sequence, the syndrome or detector pattern is computed by multiplying by the corresponding parity check matrix ($H_{X}$ or $H_{Z}$) or the graph associated with the detector error model. This binary product is performed in parallel and transmitted to the decoder under test. This is the last step of the module and has a total latency of $\left\lceil{NG/n}\right\rceil+3$ clock cycles.
2.0.2 Decoder
@@
The emulator is compatible with any quantum error correction decoder that accepts syndromes or detectors as inputs and generates errors as outputs. The interface with the rest of the emulator is straightforward; only start, ready, and done signals are needed to control the decoder. Additional information, such as the number of iterations, can be computed in the control layer or directly provided by the decoder. As an example, we have evaluated the fully parallel architecture for the scaled min-sum algorithm described in [22], applying it to multiple codes from various code families, which we will detail in the following sections.
2.0.3 Input/output parameter’s interface
@@
The emulator receives a series of parameters that can be configured at run time, avoiding the need to go through the entire synthesis and place-and-route processes. Additionally, several parameters can be queried during runtime to check how the simulation is evolving or which configuration parameters have been used. This is interesting when regions lower than $10^{-12}$ are explored, especially for long codes, when the simulations can reach daytime duration if just one board is used. The input parameters are:
@@
  • Decoder specific parameters. In the case of BP, the maximum number of iterations. It was limited to 8 bits, because of the orientation to low-latency decoders, but the number of bits associated can be modified in synthesis-time.
  • Targeted number of errors. To determine statistical significance in a simulation, designers typically set the number of Monte Carlo simulations to $100/\text{LER}$. However, because predicting the exact logical error rate can be challenging due to potential error floors, we employ a different approach. Instead of conducting a fixed number of Monte Carlo simulations, we establish a target number of erroneous patterns to identify. The simulation will continue until these targets are met.
    With 16-bit precision, each simulation can uncover up to $2^{16}$ error patterns for each physical error rate. This method allows us to generate datasets that contain a large and diverse array of error patterns. These datasets can later be analyzed offline to enhance decoding algorithms, fine-tune parameters associated with the decoder, or provide sufficient samples for training a neural network-based decoder [17].
@@
The output parameters defined in the interface are:
@@
2.0.4 Communication interface
To reduce dependence on the selected FPGA and minimize the time needed to recover information from simulations, we implemented a full stack based on Gigabit Ethernet. This includes layers for UDP, ARP, MAC, and SGMII interfaces. To enhance flexibility and protection, we designed a hardware application layer with a customized protocol. This protocol supports messages that facilitate communication with the entire emulator during and after experiments.
The primary operations supported by this protocol include starting and stopping the emulator, as well as reading and writing various parameters, such as physical error rate, logical error rate, total number of runs, average number of iterations, and clock cycles. Messages that do not conform to the protocol or are not directed to the FPGA’s MAC address are automatically filtered out and rejected.
One of the key advantages of this architecture is its capability to recover all error patterns that could not be successfully decoded. This allows for offline analysis, enabling the development of more accurate and efficient decoders.
2.0.5 Control layer

### v2

@@ Abstract
As quantum computing moves toward fault-tolerant architectures, quantum error correction (QEC) decoder performance is increasingly critical for scalability. Understanding the impact of transitioning from floating-point software to finite-precision hardware is essential, as hardware decoder performance affects code distance, qubit requirements, and connectivity between quantum and classical control units.
This paper introduces a hardware emulator to evaluate QEC decoders using real hardware instead of software models. The emulator can explore $\mathbf{10^{13}}$ different error patterns in 20 days with a single FPGA device running at 150 MHz, guaranteeing the decoder’s performance at logical rates of $\mathbf{10^{-12}}$, the requirement for most quantum algorithms. In contrast, an optimized C++ software on an Intel Core i9 with 128 GB RAM would take over a year to achieve similar results. The emulator also enables the storage of uncorrectable error patterns that generate logical errors, allowing for offline analysis and the design of new decoders.
Using results from the emulator, we propose a method that combines several belief propagation (BP) decoders with different quantization levels, which we define as a diversity-based decoder. Individually, these decoders may show subpar error correction, but together they outperform the floating-point version of BP for quantum low-density parity-check (QLDPC) codes like hypergraph or lifted product. Preliminary results with circuit-level noise and bivariate bicycle codes suggest that hardware insights can also improve software. Our diversity-based proposal achieves a similar logical error rate as the well-known approach, BP with ordered statistics (BP+OSD) decoding, with average speed improvements ranging from 30% to 80%, and 10% to 120% in worst-case scenarios, while reducing post-processing algorithm activation from 47% to 96.93%, maintaining the same accuracy.
@@ 1 Introduction
The design of decoding algorithms that obtain high accuracy and low latency for quantum low-density parity-check (QLDPC) codes [1] has been a very active area during the last decade, especially over the previous six years.
Researchers have been working on low-complexity decoding algorithms to achieve the performance of belief-propagation (BP) combined with ordered-statistics decoding (OSD) [2], [3]. This widely used approach offers good accuracy, but can be too slow for real-time decoding.
BP decoders are attractive due to their low complexity and high parallelization for real-time hardware implementations reaching low latency and allowing scalable architectures [4], [5]. However, BP introduces degradations on accuracy that are primarily due to the properties of QLDPC codes, such as degeneracy [6]. Degeneracy refers to the existence of multiple error patterns that have the same effect on the logical information, which standard BP decoding fails to properly account for. In addition, the abundance of low-weight stabilizers in QLDPC codes contributes to the formation of trapping sets, which are specific substructures in the code’s Tanner graph that do not allow messages to propagate properly through the graph. These substructures are more prevalent in QLDPC than in classical LDPC codes [7]. Trapping sets can cause the decoder to converge to incorrect solutions or oscillate between inconsistent ones, limiting the overall decoding performance. The trapping sets are one of the major contributors to the emergence of an error floor, a phenomenon where the decoder’s logical error rate (LER) stops decreasing exponentially with the improvement of the physical error rate, thus limiting accuracy.
To reduce the degradation in the LER produced by BP, OSD is used as a post-processing method, improving accuracy by several orders of magnitude [2], [3]. However, the primary drawback of OSD is that it must solve a dynamically created linear system based on the reliability information provided by BP to obtain a solution. This process of solving the linear system can create bottlenecks for implementing parallel architectures, since it involves sequential operations and dynamic computation of matrix inverses. These factors lead to scalability challenges, particularly as the size of the linear system to be solved increases, especially when considering the detector error model [8], [9].
@@
Some alternative decoders focus on avoiding OSD-like solutions that require solving systems of linear equations, which implies the computation of matrix inversion through Gaussian elimination. These alternatives aim to improve efficiency by combining different BP decoders or modifying the internal update rules of the variable and check nodes, based on the expectation that each variation enables the correction of distinct error patterns that are not successfully handled by a single BP decoder. Examples of these include Relay-BP [5], Belief Propagation Guided Decimation (BPGD) [10], [11], Stabilizer Inactivation (SI) [12], and Check-Agnosia (CA) [13].
On the other side, some decoding algorithms aim to reduce the complexity of the post-processing method, rather than using only BP decoders. Some examples are the Localized Statistics Decoder (BP+LSD) [14], the Ordered Tanner Forest (BP+OTF) [15], and Ambiguity Clustering (BP+AC) [16], which obtain high accuracy while maintaining matrix inversion to perform the decoding.
Regardless of the approach taken, all solutions begin with some variant of BP, leading to further exploration of BP algorithms to improve accuracy and reduce complexity and hence latency. This includes modifications to BP’s scheduling, such as the random reordering of node updates [17], or introducing additional noise or perturbations to address the error floor problem [18].
@@
However, several important questions remain unanswered regarding the QEC decoding algorithms. First, most quantum algorithms require a LER between $10^{-10}$ and $10^{-13}$ [19]. To ensure that, it is necessary to know the impact of finite-precision architectures on the final error correction performance of the decoders when implemented in hardware, especially in the low LER regions. To obtain statistically significant results, at least $10^{12}$ or $10^{15}$ experiments need to be run, which would be a bottleneck in software, which would require thousands of CPUs to finish the simulations within months. On the other hand, given the scaling challenges that the classical hardware will need to face to implement the QEC layer of a fault-tolerant quantum computing, it is reasonable to believe that floating-point operations may not be power or time-efficient, and accuracy should be evaluated with this limitation in mind [20].
@@
Answering these questions requires the implementation of an emulator. The scope of the emulator is to evaluate and benchmark QEC decoders in real hardware at low LERs, enabling large-scale error pattern sampling, accuracy validation, and decoder optimizations beyond the limits of software simulation. Although designing and rigorously verifying such an emulator is time-consuming, it is essential to address with warranties some of the previous problems. The accuracy of the final device will ultimately depend on the hardware architecture rather than the software implementation. The verification of the accuracy of the implemented decoders in the region of lower LERs cannot be realistically achieved through alternative methods, given limited resources. Additionally, the effects of quantization noise in finite precision implementations may lead to the development of new decoders that take advantage of the perturbations introduced by these architectures.
@@
Finally, it is important to evaluate real hardware implementations of BP, which serves as the core algorithm behind all the aforementioned state-of-the-art decoders [2], [3], [5], [10], [11], [12], [13], [14], [15], [16]. Even if other approaches, such as those based on neural networks and BP, are explored [21], utilizing a hardware emulator will enable more accurate training for achieving LERs below $10^{-12}$ (as the samples will be finite precision too), which is currently not feasible with software-based methods.
@@
  • The architecture and implementation of a hardware emulator to benchmark QEC real-time decoders and to facilitate offline analysis of the derived information. This emulator enables a significant reduction in the time required to explore low error rates down to $10^{-12}$, allowing this process to be completed in days instead of the years it would take using a software model on a CPU.
  • The design of a BP-based decoder with the knowledge obtained after studying, through the emulator, the diverse levels of noise generated by the quantization schemes. This proposal improves the LER without adding sequential post-processing algorithms like OSD. Similarly to the one in [22], in which the authors combined several noisy decoders to generate highly accurate decoding predictions. However, in our proposal, the noisy decoder is BP instead of minimum-weight perfect matching (MWPM), and the noise is not added artificially but is part of the nature of the hardware architectures, which are carefully selected. Our solution is based on just four decoders that can share the hardware in groups of two without impact on latency and a minimum memory overhead. The codes under test are hypergraph product and lifted product QLDPC codes [3], [23]. These codes were selected due to their extensive documentation in the literature under the code capacity model, which allows for verification of the emulator’s accuracy.
  • The definition of a decoder that successfully reduces both the number of BP iterations and the frequency of post-processing calls under circuit-level noise. At the same time, it maintains or improves the LER when compared to BP+OSD on bivariate bicycle codes of varying lengths and distances [24].
The structure of the paper is detailed as follows: Section II describes the proposed emulator, detailing the components of the architecture and the different configuration parameters; and introduces the verification of the platform by testing the performance of several QLDPC codes with check-node degrees of six and eight, from which some interesting conclusions are drawn. Section III presents the first diversity approach that improves the accuracy of BP decoders with moderate overhead in hardware resources. This approach leverages the quantization noise introduced by finite-precision architectures to achieve low-latency solutions. Section IV introduces a second diversity proposal based on different BP implementations, which can lead to reduced hardware requirements and shorter execution times by minimizing the number of post-processing executions. Finally, Section V concludes the paper and suggests future research lines.
@@ 2 Proposed Emulator
The architecture depicted in Fig. 1 consists of three main layers: noise and input stimulus generation (highlighted in blue in the figure), an input/output parameter interface (green), and a communication interface (orange). Additionally, there is a transversal control layer that orchestrates the entire emulator. This structure is adaptable for any type of QEC decoder (highlighted in yellow in Fig.1) that utilizes syndromes or detectors  [^1] A detector is a parity constraint on a set of measurement outcomes, as define in [8]. as inputs and produces estimated error vectors or observables as outputs. As an example, Section 2.6 analyzes the particular case of evaluating BP decoders with the emulator.
It is important to note that the implementation is vendor-agnostic, meaning it can be executed on any FPGA device, and it does not rely on proprietary IP cores  [^2] An Intellectual Property (IP) core is a reusable, modular block of hardware that implements a specific function that can be integrated into a larger hardware design to accelerate development and ensure an optimize performance when implemented on an specific technology..
@@
Next, we detail each of the modules. The code of the corresponding hardware description language of the main modules can be found in https://github.com/fpgaqec with an example for a short QLDPC code (with a 127x254 parity-check matrix). The code and architecture are fully generic, requiring only minor modifications to parameters and constants. Please note that constraint files, which map the code to hardware, are device-dependent.
2.1 Noise and input stimulus generation
@@
At the output of each of these generators, a comparator is implemented to determine whether an X, Y, or Z error has occurred at a given location. This comparator has one of the inputs wired to a programmable threshold, whose value depends on the error model and the physical error rate.
@@
Depending on the FPGA device and the length of the error vector $n$, which is equal to the codeword length for the code-capacity model, having $NG=n$ random noise generators may consume too many hardware resources, leaving insufficient space for the QEC decoder. The parameter $NG$ is generic and configurable at synthesis time. After the comparators, there is one register that receives $NG$ bits in parallel and outputs $n$ bits. So, when $NG<n$, the number of clock cycles to generate an entire error pattern is $\left\lceil{n/NG}\right\rceil$. When this error sequence is generated, it is stored in a parallel register to start producing the next noise sequence while the decoder is working. This minimizes idle periods.
@@
After the register is updated with a new sequence, the syndrome or detector pattern is computed by multiplying by the corresponding parity check matrix ($H_{X}$ or $H_{Z}$) or the graph associated with the detector error model. This binary product is performed in parallel and transmitted to the decoder under test. This is the last step of the module and has a total latency of $\left\lceil{n/NG}\right\rceil+3$ clock cycles.
Supporting both $H_{X}$ and $H_{Z}$ simulations simultaneously for realistic codes, such as the (144,12,12) QLDPC code, would require a multi-FPGA setup [5]. Even the largest FPGA devices available in 2025 (Virtex UltraScale+ V19P) can accommodate a decoder for only one type of error, with the layout nearly full, and adding real-time noise generation and the rest of the emulator parts would further increase resource demands and require precise synchronization. While the ideal setup would handle both error types concurrently, this does not affect the main findings of this work on the impact of quantization on decoding accuracy. Multi-FPGA implementations also involve substantial complexity, including multiple high-performance evaluation boards and synchronization protocols (e.g., Aurora or PCIe). We therefore note this as future work, as this study represents a first step toward a complete setup that could also consider decoders for correlated errors.
2.2 Decoder
@@
The emulator is compatible with any QEC decoder that accepts syndromes or detectors as inputs and generates errors as outputs. The interface with the rest of the emulator is straightforward; only start, ready, and done signals are needed to control the decoder. Additional information, such as the number of iterations of the decoder, can be computed in the control layer or directly provided by the decoder.
As an example, we have evaluated the fully parallel architecture for the scaled min-sum algorithm described in [4], applying it to multiple codes from various code families, which we will detail in the following sections. As a summary, the scaled min-sum algorithm is a BP-based message-passing algorithm that operates on a Tanner graph with variable and check nodes. Variable nodes represent potential error locations and combine prior information from the error model with incoming messages from check nodes, and a scaling factor $\alpha$ that is applied to the check-to-variable messages to increase accuracy. Check-nodes compute their output messages depending on whether the input messages coming from the variable nodes satisfy the syndromes or not. A hard decision (or binary decision) is made at each variable node to determine whether an error is present at that location. Decoding succeeds if this decision satisfies all syndrome constraints.
2.3 Input/output parameter’s interface
@@
The emulator receives a series of parameters to set up the noise and the accuracy of the simulation, which can be configured at run time, avoiding the need to go through the entire synthesis and place-and-route processes. Additionally, several parameters can be queried during runtime to check how the simulation is evolving or which configuration parameters have been used. This is interesting when low LERs are explored, especially for long codes, when the simulations can reach daytime duration if just one FPGA device is used. The input parameters are:
@@
  • Decoder specific parameters. In the case of BP-based decoders, the maximum number of iterations. It was limited to 8 bits, but the number of bits associated can be modified at synthesis time.
  • Targeted number of errors. To determine statistical significance in a simulation, designers typically set the number of Monte Carlo simulations to $100/\text{LER}$. However, because predicting the exact LER can be challenging due to potential error floors, we employ a different approach. Instead of performing a fixed number of Monte Carlo simulations, we stop the emulation based on a maximum number of uncorrected error patterns, i.e., 100 uncorrected error patterns. The simulation will continue until this maximum is met.
    With 16-bit precision, each simulation can uncover up to $2^{16}$ error patterns. This method allows us to generate datasets that contain a large and diverse array of error patterns. These datasets can later be analyzed offline to enhance decoding algorithms, fine-tune parameters associated with the decoder, or provide sufficient samples for training a neural network-based decoder [21].
@@
The output parameters are:
@@
2.4 Communication interface
To reduce dependence on the selected FPGA and minimize the time needed to recover information from simulations, we implemented a full stack based on Gigabit Ethernet. This includes layers for UDP, ARP, MAC, and SGMII interfaces. To enhance flexibility and provide protection against unsolicited or spurious network traffic when the FPGA operates within a shared network environment, we designed a hardware layer implementing a customized protocol to filter and control incoming messages. This protocol supports messages that facilitate communication with the entire emulator during and after experiments.
The primary operations supported by this protocol include starting and stopping the emulator, as well as reading and writing various parameters, such as physical error rate, LER, total number of runs, average number of iterations, and clock cycles. Messages that do not conform to the protocol or are not directed to the FPGA’s MAC address are automatically filtered out and rejected at the hardware level (link layer).
One important feature of this architecture is its capability to recover all error patterns that could not be successfully decoded. This allows for offline analysis, enabling the development of more accurate and efficient decoders.
2.5 Control layer